### v1

# SPHERE: Spherical partitioning for large-scale routing optimization

Abstract
We study shortest-path routing in large weighted, undirected graphs, where expanding search frontiers raise time and memory costs for exact solvers. We propose SPHERE, a source-target-aware heuristic that identifies an $s$-$t$ overlap: vertices that are close to both $s$ and $t$ in hop count. Selecting an anchor $a$ in this overlap partitions the task into two subproblems with unchanged problem-topology , $s\to a$ and $a\to t$; if either remains large, the procedure recurses on its induced subgraph. Because the cut lies inside the overlap, concatenating the resulting subpaths yields a valid $s\to t$ route without boundary repair. SPHERE is independent of the downstream solver (e.g., Dijkstra) and exposes parallelism across subproblems. On large networks, it achieves faster runtimes and smaller optimality gaps than Louvain-based routing and a METIS-based pipeline, even on graphs with more than a million nodes and edges, while also outperforming Dijkstra in runtime.
1 Introduction
Point-to-point routing in large weighted, undirected graphs is a canonical problem in operations research, one that emerged alongside the field’s inception.. Exact label-setting or label-correcting methods (e.g., Dijkstra) can certify optimality (Dijkstra, 1959), but their search frontiers grow quickly on large networks and become memory- and time-intensive. Graph partitioning methods (e.g., Kernighan-Lin, multilevel partitioning/METIS, spectral cuts) reduce problem size by cutting the graph into parts (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Shi and Malik, 2000), yet cuts chosen without regard to a specific query $(s,t)$ can intersect narrow passages that many feasible sstt routes must use, which can exclude high-quality routes or require costly boundary repair. Bidirectional search reduces depth by growing frontiers from $s$ and $t$ (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997; Holte et al., 2016), but still explores a single coupled state space and may expand many partial routes on large graphs. For completeness, we note that while feasibility for unconstrained shortest paths is trivial on connected graphs, constrained variants (e.g., resource- or budget-limited paths) are NP-hard; our focus here is the unconstrained case with nonnegative weights, which admits an optimal solution via Dijkstra’s algorithm (Dijkstra, 1959).
We propose SPHERE, a query-aware partition that aligns computation to the sstt geometry. Let $B_{s}(r)$ and $B_{t}(r)$ denote hop-distance spheres around $s$ and $t$. SPHERE starts from a radius tuple $(r_{s},r_{t})$ with guaranteed overlap $B_{s}(r_{s})\cap B_{t}(r_{t})\neq\emptyset$, then applies a monotone decrement rule that reduces the radii (when needed, both simultanously) so that $|r_{s}-r_{t}|\leq 1$ while preserving overlap, and selects the minimal balanced pair, i.e. no further admissible unit decrement of either radius (still keeping $|r_{s}-r_{t}|\leq 1$) leaves the intersection nonempty. Then it selects an anchor $a\in B_{s}(r_{s})\cap B_{t}(r_{t})$ and splits the query into $s\!\to\!a$ and $a\!\to\!t$, solved on the corresponding induced subgraphs under a tunable size cap. If a subproblem still exceeds the cap, the procedure recurses on its induced subgraph. Because the cut lies inside the overlap, concatenating subpaths at $a$ yields a valid sstt route without boundary repair. The partition is solver-agnostic and the independent subproblems can be solved in parallel.
Our contributions are:
  1. 1.
    Query-aware partition. We introduce a simple heuristic that adaptively splits the graph based on the geometry of the source-target pair, ensuring that each part preserves connectivity and can be rejoined without additional repair.
  2. 2.
    Parallel subinstances. The subproblems generated by SPHERE are independent induced subgraphs, naturally exposing parallelism and enabling scalable routing on distributed systems with effort scaling to hardware.
  3. 3.
    Empirical validation. On large road graphs, SPHERE achieves faster runtimes than Louvain-based routing and a METIS-based pipeline while matching or improving route quality, benefiting further from inherent parallelism.
SPHERE is complementary to downstream solvers and can wrap Dijkstra or A* (Dijkstra, 1959; Hart et al., 1968), as well as modern speedup indices such as ALT/landmarks (Goldberg and Harrelson, 2005), reach-based pruning (Gutman, 2004), contraction hierarchies (Geisberger et al., 2008), hub labeling (Abraham et al., 2011), and transit-node routing (Bast et al., 2016); we evaluate with common solvers and baselines and discuss integration in Section 5. Next, we motivate these design choices by examining where existing approaches trade-off feasibility, and instance size, before positioning SPHERE within the literature and formalizing our approach.
2 Motivation
Routing at scale must meet three criteria simultaneously: control per-instance complexity, make an $s$-$t$ route feasible, and remain compatible with trusted downstream solvers while enabling parallel execution. Existing tools typically sacrifice at least one of these: Labeling methods (Feillet et al., 2004; Irnich and Desaulniers, 2005) preserve feasibility but generate a combinatorial explosion of partial routes on large graphs. Graph partitioners (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Shi and Malik, 2000) control instance size, but because their boundaries are chosen without regard to $s$ and $t$, they may cut across a bottleneck that every feasible sstt path must traverse, risking the loss of valid routes. Bidirectional search (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997; Holte et al., 2016) reduces search depth by expanding from both ends, $s$ and $t$, but the two frontiers still form a single combined search space and therefore cannot be partitioned into smaller, independent pieces suitable for parallel solution.
SPHERE overcomes these trade-offs by aligning the partition with the specific request. To ensure controllable size, it expands hop-distance spheres from $s$ and $t$ until reaching the last nonempty overlap, where a user-defined radius limits the diameter of each subproblem. Second, for feasibility, SPHERE selects an anchor $a$ in the overlap and solves $s\!\to\!a$ and $a\!\to\!t$ on their respective induced subgraphs, so cuts always lie inside the overlap and concatenation at $a$ is guaranteed to generate a feasible sts-t route. This preserves all local connectivity within each part and ensures that every feasible subpath remains feasible in the global graph. Third, the resulting subproblems are independent, enabling parallel execution across multiple cores, and the design is solver-agnostic which means that any downstream optimizer, from exact Mixed-Integer Programming (MIP) on small pieces to strong metaheuristics on larger ones (Pisinger and Ropke, 2007; Vidal, 2022), can be applied without modification. In short, SPHERE offers a partition strategy that controls complexity, preserves feasibility, and exposes natural parallelism while remaining compatible with the solvers practitioners already trust.
3 Related Work
Point-to-point routing on large weighted graphs has been studied from several perspectives. Classical exact methods such as label-setting and label-correcting algorithms (e.g., Dijkstra) can guarantee global optimality on connected graphs with nonnegative weights (Dijkstra, 1959). Dynamic-programming and labeling variants extend this to richer settings and can certify optimality, but they often face an explosion in the number of labels on large graphs (Feillet et al., 2004; Irnich and Desaulniers, 2005). Approximation schemes exist for restricted variants (Hassin, 1992), trading accuracy for speed, yet these techniques operate on the full graph and therefore do not reduce the instance size before optimization.
Graph partitioning offers a complementary idea: simplify the problem by splitting the graph. Kernighan-Lin improves an initial cut by greedy node swaps to reduce the number of edges crossing between parts while keeping parts balanced (Kernighan and Lin, 1970). Multilevel schemes such as METIS repeatedly coarsen the graph, compute a cut on the coarse instance, and refine it during uncoarsening, which gives strong scalability in practice (Karypis and Kumar, 1998). In routing pipelines, these partitions are often used to build a quotient graph for a coarse path that is later refined within parts. A limitation of all such methods is that the cut is chosen without regard to the specific start-target pair $(s,t)$, so the boundary can lie across a narrow passage that every feasible $s$-$t$ path must use. Once this happens, the coarse or restricted routing stage may miss or degrade the necessary connection.
Bidirectional search reduces search depth by expanding frontiers from both ends and meeting near the middle (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997). The MM algorithm of Holte et al. (2016) provides a priority rule that ensures expansions do not go beyond the midpoint of the optimal solution. These methods improve exploration efficiency but still operate within a single coupled search space and do not produce small, independent instances. In large graphs this can still generate many partial routes, and unlike partition-based methods, they do not offer a straightforward path to parallel, solver-agnostic execution.
4 Spherical Partitioning
4.1 Spherical subgraphs
We begin by fixing distance and the local “window” around a node. In the following, let every graph $G=(V,E)$ be a finite, connected, undirected graph with positive edge weights. Unless stated otherwise we measure distance in hops (unweighted shortest path); all statements below remain valid for the weighted shortest-path distance.
Definition 1 (Spheres).
For $R\in\mathbb{N}$ and a reference node $\tilde{v}\in V$, the (closed) sphere is
$S_{R}(\tilde{v})\;:=\;\{\,v\in V:\ d_{G}(\tilde{v},v)\leq R\,\}\subseteq V,$
where $d_{G}(\tilde{v},v)$ is the unweighted shortest path distance between $\tilde{v}$ and $v$.
[…]
Definition 2 (Spherical subgraph).
[…]
$E^{\mathrm{sp}}_{R}(\tilde{v})=\bigcup_{v\in S_{R}(\tilde{v})}\mathcal{E}\!\bigl(\mathrm{SP}(\tilde{v},v)\bigr),$
[…]
Definition 3 (Induced spherical subgraph).
[…]
4.2 Partition cut
With these building blocks in place, we now align the local windows to a specific query $(s,t)$. For radii $(R_{s},R_{t})\in\mathbb{N}^{2}$, define the overlap
$O(R_{s},R_{t})\;:=\;S_{R_{s}}(s)\cap S_{R_{t}}(t).$
We control how radii evolve, how they are initialized, and how an anchor is chosen in the overlap.
Definition 4 (Decrement rule).
A decrement rule is a map $\mathrm{decru}:\mathbb{N}^{2}\to\mathbb{N}^{2}$ that is monotone in the lexicographic order, i.e.,
$(x_{1},x_{2})\leq_{\mathrm{lex}}(y_{1},y_{2})\;\Longleftrightarrow$ $\begin{aligned} &x_{1}<y_{1},\\ &\text{or }(x_{1}=y_{1}\text{ and }x_{2}\leq y_{2}).\end{aligned}$
and never increases any coordinate.
Example (monotone decrement).
A simple rule that satisfies the definition is
$\mathrm{decru}(R_{s},R_{t})=\begin{cases}(R_{s}{-}1,R_{t}),&R_{s}\geq R_{t},\\[2.0pt] (R_{s},R_{t}{-}1),&\text{otherwise.}\end{cases}$
It is coordinatewise nonincreasing, and it is lex-monotone: if $(R_{s},R_{t})\leq_{\mathrm{lex}}(R^{\prime}_{s},R^{\prime}_{t})$, then either $R_{s}<R^{\prime}_{s}$ (so the first branch cannot produce a larger first coordinate) or $R_{s}=R^{\prime}_{s}$ with $R_{t}\leq R^{\prime}_{t}$, in which case both images either decrement the same coordinate or the smaller pair cannot “overtake” the larger one in lexicographic order.
The map $\mathrm{decru}$ from Definition 4 determines how radii shrink once chosen. To make the procedure operative, we also need: (i) an initial choice of radii that guarantees the spheres overlap at the start, and (ii) a way to convert any nonempty overlap into a concrete split point. These are captured by the following two rules.
Definition 5 (Starting rule).
A starting rule $\mathrm{staru}:V\times V\to\mathbb{N}^{2}$ returns $(R_{s}^{0},R_{t}^{0})$ with nonempty overlap $O(R_{s}^{0},R_{t}^{0})\neq\emptyset$ (e.g., $R_{s}^{0}{+}R_{t}^{0}\geq d_{G}(s,t)$; a stronger but safe choice is $S_{R_{s}^{0}}(s)\cup S_{R_{t}^{0}}(t)=V$).
Definition 6 (Anchor rule).
An anchor rule $\mathrm{anru}:\mathcal{P}(V)\to V$ selects $a\in O(R_{s},R_{t})$ (e.g., uniform random).
Bringing these pieces together, the three rules: starting rule for initialization (Definition 5), decrement rule for shrinking (Definition 4), and anchor rule for selecting the split point (Definition 6), jointly define the following algorithm, which we refer to as the Partition cut:
Algorithm 1 Partition cut
1:graph $G=(V,E)$, terminals $(s,t)$,
2:     rules $(\mathrm{staru},\mathrm{decru},\mathrm{anru})$
3:  radii $(\bar{R}_{s},\bar{R}_{t})$, overlap $O$, anchor $a$
[…]
5:while $O(R_{s},R_{t})\neq\emptyset$ do
6:  $(\bar{R}_{s},\bar{R}_{t})\leftarrow(R_{s},R_{t})$
7:  $(R_{s},R_{t})\leftarrow\mathrm{decru}(R_{s},R_{t})$
8:end while
[…]
10:return $(\bar{R}_{s},\bar{R}_{t},O,a)$
The intuition behind Algorithm 1 can be explained as follows and is visualized in Figure 2. Grow spheres around $s$ and $t$ until they overlap, then shrink via $\mathrm{decru}$ until the next shrink would empty the overlap. Pick an anchor $a$ according to Definition 6 in this last nonempty overlap, and split the query into $s\!\to\!a$ and $a\!\to\!t$ subproblems on the corresponding induced spheres.
[…]
Figure 2: Last nonempty overlap. Blue and red spheres overlap in the green region containing both $a$ and $a^{\prime}$. Potential anchor node $a$ is selected according to Definition 6.
A natural question is whether the Partition cut of Algorithm 1 actually preserves feasibility of the original $s{\to}t$ routing problem. In other words, does splicing the two subpaths at the chosen anchor still yield a valid global route? The following statement makes this precise:
Proposition 4.1 (Feasibility preservation).
Let $(\bar{R}_{s},\bar{R}_{t},O,a)$ be returned by Algorithm 1. If $P_{s\to a}$ is a feasible path in $G^{\mathrm{ind}}_{\bar{R}_{s}}(s)$ and $P_{a\to t}$ is a feasible path in $G^{\mathrm{ind}}_{\bar{R}_{t}}(t)$, then $P_{s\to a}\!\circ\!P_{a\to t}$ is a feasible $s\to t$ path in $G$.
Proof sketch.
Both subgraphs are induced, hence every used edge is in $G$. Since $a\in O(\bar{R}_{s},\bar{R}_{t})$, concatenating at $a$ yields a valid $s\to t$ path in $G$. ∎
4.3 Recursive spherical partitioning
The following algorithm extends the partition cut from Algorithm 1 into a recursive partitioning routine. Its goal is to partition a global $s\!\to\!t$ routing task into a set of smaller subproblems, each confined to an induced sphere whose radius does not exceed a user-defined bound $R_{\max}$. These subproblems are returned as task triples $(H,u,w)$, meaning: “route from $u$ to $w$ within the induced subgraph HH.”
Algorithm 2 Recursive spherical partitioning
1graph $G=(V,E)$, terminals $(s,t)$,
2     radius cap $R_{\max}$, rules $\mathrm{staru},\mathrm{decru},\mathrm{anru}$
3  set $\mathcal{T}$ of task triples $(H,u,w)$
4function SphPartition($G,s,t,R_{\max}$)
5  $\mathcal{T}\leftarrow\emptyset$, $\texttt{rule}\leftarrow(\mathrm{staru},\mathrm{decru},\mathrm{anru})$
6  $(R_{s},R_{t},O,a)\leftarrow\textsc{PartCut}(G,s,t,\texttt{rule})$ $\triangleright$ Alg. 1
7  if $R_{s}\leq R_{\max}$ then
8  $\mathcal{T}\leftarrow\mathcal{T}\cup\{(G^{\mathrm{ind}}_{R_{s}}(s),\,s,\,a)\}$
9  else
10  $\mathcal{T}\leftarrow\mathcal{T}\cup\textsc{SphPartition}(G^{\mathrm{ind}}_{R_{s}}(s),\,s,\,a,\,R_{\max})$
11  end if
12  if $R_{t}\leq R_{\max}$ then
13  $\mathcal{T}\leftarrow\mathcal{T}\cup\{(G^{\mathrm{ind}}_{R_{t}}(t),\,a,\,t)\}$
14  else
15  $\mathcal{T}\leftarrow\mathcal{T}\cup\textsc{SphPartition}(G^{\mathrm{ind}}_{R_{t}}(t),\,a,\,t,\,R_{\max})$
16  end if
17  return $\mathcal{T}$
18end function
Algorithm 2 starts by initializing the set of tasks $\mathcal{T}$ to be empty (lines 1-3). It then performs a partition cut on the original problem $(G,s,t)$, which returns the radii $(R_{s},R_{t})$, the last nonempty overlap $O$, and an anchor $a\in O$. The anchor acts as a splitting point that separates the global $s\!\to\!t$ route into two sides.
The algorithm then branches into two recursive checks. On the $s$-side (lines 4-7), the relevant subproblem is $(G^{\mathrm{ind}}_{R_{s}}(s),s,a)$. If the radius $R_{s}$ is already below the cap $R_{\max}$, this subproblem is considered sufficiently small and is added to the task set $\mathcal{T}$. If not, the procedure calls itself recursively on this induced subgraph to split it further. Symmetrically, the $t$-side branch (lines 8-11) processes the subproblem $(G^{\mathrm{ind}}_{R_{t}}(t),a,t)$: either it is recorded directly if $R_{t}\leq R_{\max}$, or it is further partitioned by another recursive call.
Finally, once both branches have been processed, the algorithm returns (line 12) the accumulated set $\mathcal{T}$ of task triples. By construction, concatenating the solutions to these recorded subproblems along their anchors yields a valid overall $s\!\to\!t$ route.
(G,s,t)(G,s,t)Partition cut $\Rightarrow$$(R_{s},R_{t},O,a)$$G^{\mathrm{ind}}_{R_{s}}(s),\ s\to a$$G^{\mathrm{ind}}_{R_{t}}(t),\ a\to t$if $R_{s}\leq R_{\max}$: record; else: recurseif $R_{t}\leq R_{\max}$: record; else: recurse
Figure 3: Recursive splitting. Any side above the radius cap is split again by Algorithm 1; leaves are independent subproblems.
Figure 3 illustrates this recursive structure. Each partition cut splits the problem into two boxes, one for the start-side and one for the target-side. If the radius on a side is within the threshold, the box is recorded as a leaf; otherwise, it is expanded by another partition cut.
4.3.1 Baseline application
The recursive partition in Algorithm 2 produces a set of leaf tasks, each represented by a triple $(H,u,w)$ where $H$ is an induced spherical subgraph and $u,w$ are its boundary terminals. At this stage, the global $s\to t$ routing problem has been reduced to a collection of independent local subproblems, each small enough to be solved efficiently by standard methods.
Algorithm 3 Baseline application with spherical partitioning
1:graph $G=(V,E)$, terminals $(s,t)$,
2:     radius cap $R_{\max}$, solver $\mathrm{optimize}(\cdot)$
3:  feasible route $\mathcal{R}$
[…]
5:for all $(H,u,w)\in\mathcal{T}$ do
6:  $\mathcal{R}\leftarrow\mathcal{R}\circ\mathrm{optimize}(H,u,w)$
7:end for
8:return $\mathcal{R}$
Algorithm 3 describes how to turn the recursive spherical partitioning into a full routing solution. First, the input graph $G$ with terminals $s,t$ is partitioned into a collection of smaller subproblems (line 3). Each subproblem $(H,u,w)$ is an induced subgraph $H$ together with its local terminals $u$ and $w$. Then, for each subproblem, an arbitrary downstream solver $\mathrm{optimize}(H,u,w)$ is applied to compute a local path (lines 4-5). These local solutions are concatenated in the order determined by the anchors of the partition, resulting in a valid global route $\mathcal{R}$ from $s$ to $t$ (line 6). Since the subproblems are independent, they can also be solved in parallel, and the procedure remains agnostic to the choice of solver.
5 Numerical Experiments
Our experiments are designed to evaluate whether our method delivers significant performance gains (in terms of runtime and optimality gap) over established graph-routing baselines.
5.1 Dataset
For our empirical evaluation we use the well-known West-USA-road benchmark suite from the 9th DIMACS Implementation Challenge on Shortest Paths (DIM, 2006). These graphs encode real-world road networks at the level of individual intersections and road segments. Each node corresponds to a road intersection or endpoint, and each edge represents a street segment with a positive weight equal to its physical length. The dataset is widely used in the routing literature as a benchmark for large-scale shortest-path computations, since it combines realistic geometry, heterogeneous edge weights, and large graph sizes (millions of nodes and edges).
5.2 Experimental setup
For each problem seed $p\in\{1,\dots,30\}$ we draw a source-target pair $(s,t)$ uniformly at random from the node set (rejecting $s=t$). For each $(s,t)$, we run five independent trials with inner seeds $q\in\{1,\dots,5\}$ to capture variability from randomized components. For instance, the anchors are selected randomly according to Definition 6. We report results per instance and aggregate by averaging across the $5$ inner seeds for each problem seed.
Given $(s,t)$, we apply Algorithm 2 and selects an anchor node uniformly at random from this set of nodes (the randomness is controlled by the inner seed $q$). The graph is partitioned into subgraphs around $s$ and $t$ until a maximum radius threshold is met. $R_{max}$ from Algortihm 3 is set to $1800$. Throughout development we found a robust rule of thumb: when routing within a spherical subgraph of hop radius $R$, choosing $R_{\max}\in[\,R/2,\;R/1.7\,]$ delivers high-quality solutions at both low computational cost and effort. For the West-USA instances, a single sphere with $R\approx 3200$ covers the region; thus $R_{\max}=1800$ lies squarely in this band, balancing overlap likelihood and leaf size. Each subgraph shortest-path subproblem is solved with Dijkstra’s algorithm. Because our spherical partitioning produces independent induced subgraphs, we parallelize these subgraph solves and report end-to-end wall-clock time (including partition cut and final concatenation). By contrast, the baselines do not naturally partition into independent leaf tasks so we do not apply the same solving subproblem parallelization to them.
Baselines.
As baselines, we first include an optimal reference obtained by running Dijkstra’s algorithm directly on the full graph. To compare against partition-based approaches, we also evaluate a corridor-routing strategy in which the graph is partitioned into $k=64$ cells using PyMetis; a path is computed on the resulting quotient graph, and Dijkstra’s algorithm is then applied within the induced corridor to recover a detailed route (Karypis and Kumar, 1998). In addition, we test a partition heuristic based on community detection via the Louvain method: routing is first performed on the community graph, and the solution is subsequently refined by applying Dijkstra on the induced subgraph spanned by the relevant communities (default parameters; seeded where applicable) (Blondel et al., 2008). The corridor-routing and Louvain pipelines introduce partitions, yet can not be solved independently and not conditioned on the specific $(s,t)$, so they likewise do not yield independent leaves that one could solve with a plug-in solver and then concatenate repair-free (e.g. fixing feasibility) (Karypis and Kumar, 1998; Blondel et al., 2008).
Metrics. Let $m\in\{\textsc{Sphere},\ \textsc{Corridor},\ \textsc{Louvain}\}$ denote the method. For each problem seed $p$ and inner seed $q$, let $C_{p,q}^{(m)}$ be the route cost returned by $m$, and let $C_{p}^{\star}$ be the shortest-path cost from Dijkstra’s algorithm on the full graph for the same $(s,t)$. Define the relative optimality gap
$\Delta_{p,q}^{(m)}\;=\;\frac{C_{p,q}^{(m)}-C_{p}^{\star}}{C_{p}^{\star}}.$
For each $p$, aggregate across the five inner seeds by the median:
$\widetilde{\Delta}_{p}^{(m)}\;=\;\operatorname{median}_{q=1,\dots,5}\,\Delta_{p,q}^{(m)},$
$\widetilde{T}_{p}^{(m)}\;=\;\operatorname{median}_{q=1,\dots,5}\,T_{p,q}^{(m)},$
where $T_{p,q}^{(m)}$ is the end-to-end wall-clock time. For completeness, we also report means:
$\overline{\Delta}_{p}^{(m)}\;=\;\tfrac{1}{5}\sum_{q=1}^{5}\Delta_{p,q}^{(m)},\qquad\overline{T}_{p}^{(m)}\;=\;\tfrac{1}{5}\sum_{q=1}^{5}T_{p,q}^{(m)}.$
5.3 Results
Across 30 West-USA instances with five inner seeds each, SPHERE achieves faster end-to-end runtimes than the three baselines and matches near-optimal route quality. We use performance and accuracy profiles for visualization (Dolan and Moré, 2002; Beiranvand et al., 2017). Both are computed per instance from the median over seeds. Mean-based profiles for runtime and accuracy are presented as well.
(a) Runtime profile based on the median over five seeds
(b) Accuracy profile based on the median over five seeds
(c) Runtime profile based on the mean over five seeds
(d) Accuracy profile based on the mean over five seeds
Figure 4: Performance and accuracy profiles of 30 instances based on median and mean over five seeds.
Runtime.
Figures 4(a) and 4(c) plot the fraction of instances solved within a performance ratio $\tau$ of the best method on that instance. SPHERE’s curve lies above the Louvain and corridor routes for most $\tau$, indicating uniformly stronger or equal runtime across instances. Interestingly, for small $\tau$, Louvain’s curve lies below Dijkstra’s, despite being a suboptimal method. However, at larger $\tau$ it crosses and stays above, indicating tail slowdowns for Dijkstra. This suggests that SPHERE not only improves median runtime, but also avoids the heavy-tailed slowdowns observed in both Dijkstra and Louvain, resulting in tighter runtime distributions and fewer extreme outliers, which makes SPHERE’s runtimes more predictable across different instances.
Route quality.
Figures 4(b) and 4(d) reports the fraction of instances whose relative optimality gap is at most $\tau$. Dijkstra attains zero gap by construction. SPHERE’s curve dominates the partition baselines across $\tau$, yielding smaller or equal gaps on the majority of instances. This confirms that cutting inside the $s$-$t$ overlap and solving on induced spheres preserves high-quality routes without boundary repair.
Pareto dominance
Against the partition baselines (Louvain and METIS corridor), SPHERE Pareto-dominates on 26 of 30 instances in terms of average metrics. It is both faster and has a smaller optimality gap (Appendix A.3). The remaining three show a speed-quality trade-off where SPHERE is not dominated by another algorithm. In other words, SPHERE consistently lies on or close to the empirical speed–quality frontier, while Louvain and corridor-routing often face a trade-off between being faster but less accurate, or more accurate but substantially slower.
Robustness and aggregation
Repeating the analysis with means over seeds yields profiles that do not differ materially from the median-based plots, supporting robustness to the randomized anchor selection. Additionally, the full results in the appendix show low dispersion across seeds. For SPHERE, the optimality-gap standard deviation is 0 on 16 instances and is the lowest among all methods on all but one instance. For further Information, we refer to Appendix A.3.
6 Conclusion and Outlook
We introduced SPHERE, a routing scheme that splits an sstt query by cutting inside the last nonempty overlap of hop-spheres around $s$ and $t$. This placement preserves feasibility without boundary repair, confines each side to an induced subgraph, and allows the two subproblems to be solved independently with standard shortest-path routines and in parallel.
On West-USA-road graphs, SPHERE achieves faster end-to-end times and equal or smaller optimality gaps than a Louvain-based route and a METIS-corridor pipeline on the large majority of instances; runtimes are also more predictable, as heavy-tail slowdowns seen for the baselines largely disappear. In contrast to bidirectional search, which still explores a single coupled state space, SPHERE yields separable leaves. In contrast to static partitions whose cuts ignore $(s,t)$ and can sever bottlenecks, cutting within the overlap keeps the route intact and avoids corridor repair.
Looking ahead, the main levers are the starting rule, the decrement rule, and the anchor selection. Even simple choices already perform well, but data-driven or adaptive variants should improve robustness across graphs and query distributions. A second priority is to tune the radius cap $R_{\max}$ automatically (per instance or region) to balance overlap likelihood against subproblem size. These steps retain the current simplicity and parallelism while targeting further speed and stability.
Appendix A Appendix
A.1 Notation
{longtblr}
[ caption = Notation, label = tab:notation ]colspec = @ Q[l,wd=4.6cm] X @, rowhead = 2 Symbol Description
$V=\{1,\dots,m\}$ Node (vertex) index set.
$\mathcal{P}(V)$ Power set of $V$.
$E\subseteq V\times V$ Edge set; undirected unless stated.
$G=(V,E)$ Graph with node set $V$ and edge set $E$.
$d_{G}(u,v)$ Shortest unweighted path distance on $G$.
$G[U]$ Induced subgraph on node set $U$.
$S_{R}(\tilde{v})=\{\,v\in V:\ d_{G}(\tilde{v},v)\leq R\,\}$ (Closed) hop sphere (ball) of radius $R$ around $\tilde{v}$.
$S_{s}(r)\equiv S_{r}(s),\;S_{t}(r)\equiv S_{r}(t)$ Shorthand spheres around $s$ and $t$.
$O(R_{s},R_{t})=S_{R_{s}}(s)\cap S_{R_{t}}(t)$ Overlap of the two spheres.
$a\in V$,  $S^{\dagger}\subseteq O(R_{s},R_{t})$ Anchor node; optional candidate anchor set.
$G_{s}=G[S_{R_{s}}(s)],\;G_{t}=G[S_{R_{t}}(t)]$ Induced subgraphs around $s$ and $t$ at radii $R_{s},R_{t}$.
$(H,u,w)$,  $\mathcal{T}$ Task triple “route $u\to w$ inside HH”; set of leaf tasks.
$|X|$ Cardinality of a (finite) set $X$.
A.2 Parameters
{longtblr}
[ caption = Parameters, label = tab:params ]colspec = @ Q[l,wd=3.4cm] X @, rowhead = 2 Symbol Description
$R_{\max}\in\mathbb{N}$ Maximum allowed radius for spherical subgraphs (controls subproblem diameter).
$V_{\max}\in\mathbb{N}$,  $E_{\max}\in\mathbb{N}$ Optional caps on nodes/edges permitted in any subgraph $G[U]$.
$\varepsilon_{\text{overlap}}\in\mathbb{N}_{0}$ Minimum required overlap size (require $|O(R_{s},R_{t})|\geq\varepsilon_{\text{overlap}}$).
$\Delta r\in\mathbb{N}$ Radius increment/decrement step during sphere growth/shrinking (if using fixed steps; typically $\Delta r=1$).
$L_{\max}\in\mathbb{N}$ Maximum recursion depth for repeated splits.
$k_{\text{anchor}}\in\mathbb{N}$ Number of anchors to select from $O$ (often $k_{\text{anchor}}=1$).
A.3 Performance per instance
Table 1: WestUSA – Routing
avg compute time avg opt. gap median opt. gap std opt. gap
P1 SPHERE 135.73 0.08 0.06 0.02
Corridor Route 173.97 0.24 0.07 0.36
Louvain 1047.77 0.25 0.28 0.05
P2 SPHERE 138.27 0.12 0.12 0.00
Corridor Route 187.87 0.14 0.09 0.16
Louvain 1072.17 0.34 0.35 0.05
P3 SPHERE 165.25 0.03 0.03 0.00
Corridor Route 190.12 0.08 0.10 0.06
Louvain 1130.18 0.18 0.18 0.05
P4 SPHERE 370.06 0.11 0.11 0.00
Corridor Route 574.18 0.25 0.26 0.10
Louvain 1230.71 0.37 0.37 0.03
P5 SPHERE 126.88 0.05 0.05 0.00
Corridor Route 171.35 0.10 0.11 0.06
Louvain 1129.89 0.14 0.14 0.06
P6 SPHERE 301.07 0.02 0.03 0.00
Corridor Route 571.56 0.18 0.15 0.14
Louvain 1265.49 0.25 0.20 0.12
P7 SPHERE 213.59 0.00 0.00 0.00
Corridor Route 354.04 0.27 0.36 0.17
Louvain 1197.91 0.32 0.32 0.04
P8 SPHERE 157.02 0.01 0.01 0.00
Corridor Route 224.89 0.32 0.27 0.16
Louvain 1165.53 0.09 0.07 0.03
P9 SPHERE 390.68 0.02 0.02 0.00
Corridor Route 396.99 0.13 0.10 0.10
Louvain 1274.68 0.23 0.21 0.06
P10 SPHERE 137.20 0.02 0.02 0.00
Corridor Route 197.31 0.02 0.00 0.03
Louvain 1141.72 0.27 0.33 0.10
P11 SPHERE 130.96 0.13 0.13 0.01
Corridor Route 167.76 0.20 0.07 0.24
Louvain 1057.36 0.33 0.33 0.00
P12 SPHERE 264.38 0.04 0.04 0.00
Corridor Route 234.53 0.24 0.20 0.13
Louvain 1160.17 0.25 0.17 0.13
P13 SPHERE 123.22 0.00 0.00 0.00
Corridor Route 200.92 0.10 0.00 0.16
Louvain 1308.92 0.18 0.19 0.03
P14 SPHERE 234.34 0.14 0.14 0.00
Corridor Route 358.43 0.25 0.24 0.17
Louvain 1414.78 0.33 0.31 0.05
P15 SPHERE 211.37 0.00 0.00 0.00
Corridor Route 300.88 0.11 0.10 0.09
Louvain 1288.33 0.15 0.15 0.02
P16 SPHERE 136.40 0.04 0.04 0.00
Corridor Route 221.84 0.08 0.02 0.13
Louvain 1400.16 0.20 0.19 0.05
P17 SPHERE 119.90 0.01 0.00 0.02
Corridor Route 193.90 0.02 0.00 0.03
Louvain 1266.58 0.20 0.21 0.07
P18 SPHERE 123.18 0.08 0.08 0.01
Corridor Route 205.16 0.06 0.00 0.07
Louvain 1409.96 0.32 0.33 0.03
P19 SPHERE 205.97 0.02 0.02 0.00
Corridor Route 1052.15 0.40 0.30 0.31
Louvain 1401.55 0.25 0.25 0.05
P20 SPHERE 170.12 0.00 0.00 0.00
Corridor Route 391.16 0.30 0.08 0.34
Louvain 1414.17 0.10 0.08 0.04
P21 SPHERE 326.25 0.01 0.01 0.00
Corridor Route 506.65 0.20 0.14 0.12
Louvain 1502.53 0.30 0.22 0.16
P22 SPHERE 137.27 0.03 0.03 0.00
Corridor Route 180.60 0.49 0.29 0.51
Louvain 1135.72 0.28 0.26 0.08
P23 SPHERE 172.50 0.01 0.01 0.00
Corridor Route 211.28 0.08 0.04 0.11
Louvain 1065.85 0.21 0.20 0.06
P24 SPHERE 145.29 0.02 0.02 0.00
Corridor Route 202.11 0.21 0.26 0.13
Louvain 1009.77 0.27 0.28 0.07
P25 SPHERE 160.44 0.02 0.02 0.00
Corridor Route 219.89 0.11 0.04 0.14
Louvain 1059.09 0.12 0.12 0.02
P26 SPHERE 172.71 0.03 0.03 0.00
Corridor Route 203.42 0.17 0.13 0.14
Louvain 1054.63 0.24 0.19 0.13
P27 SPHERE 246.94 0.02 0.02 0.00
Corridor Route 479.93 0.14 0.06 0.11
Louvain 1388.28 0.22 0.20 0.04
P28 SPHERE 159.42 0.00 0.00 0.00
Corridor Route 267.58 0.16 0.13 0.17
Louvain 1274.86 0.19 0.16 0.08
P29 SPHERE 232.46 0.07 0.07 0.00
Corridor Route 354.39 0.14 0.11 0.07
Louvain 1219.90 0.25 0.24 0.03
P30 SPHERE 142.64 0.16 0.16 0.00
Corridor Route 175.93 0.10 0.08 0.10
Louvain 1162.54 0.15 0.09 0.10

### v2

@@
 We study shortest-path routing in large weighted, undirected graphs, where expanding search frontiers raise time and memory costs for exact solvers.
-We propose SPHERE, a source-target-aware heuristic that identifies an $s$-$t$ overlap: vertices that are close to both $s$ and $t$ in hop count.
-Selecting an anchor $a$ in this overlap partitions the task into two subproblems with unchanged problem-topology , $s\to a$ and $a\to t$; if either remains large, the procedure recurses on its induced subgraph.
-Because the cut lies inside the overlap, concatenating the resulting subpaths yields a valid $s\to t$ route without boundary repair.
-SPHERE is independent of the downstream solver (e.g., Dijkstra) and exposes parallelism across subproblems.
-On large networks, it achieves faster runtimes and smaller optimality gaps than Louvain-based routing and a METIS-based pipeline, even on graphs with more than a million nodes and edges, while also outperforming Dijkstra in runtime.
+We propose SPHERE, a query-aware partitioning heuristic that adaptively splits the problem by identifying source-target (sstt) overlaps of hop-distance spheres.
+Selecting an anchor node $a$ within this overlap partitions the task into independent induced subgraphs for $s\to a$ and $a\to t$, each restricted to its own induced subgraph.
+If resulting subgraphs remain large, the procedure recurses on that specific subgraph.
+We provide a formal guarantee that by using the partition cut within the shared overlap, the resulting subpaths preserve feasibility, thereby avoiding the need for boundary repair.
+Furthermore, SPHERE acts as a solver-agnostic framework that naturally exposes parallelism across subproblems.
+On million-scale road networks, SPHERE achieves faster runtimes and smaller optimality gaps than contemporary state-of-the-art partitioning and community-based routing pipelines.
+Crucially, it also substantially mitigates heavy-tail runtime outliers suffered by standard exact methods, yielding highly stable and predictable execution times across varying queries.
 1 Introduction
-Point-to-point routing in large weighted, undirected graphs is a canonical problem in operations research, one that emerged alongside the field’s inception..
-Exact label-setting or label-correcting methods (e.g., Dijkstra) can certify optimality (Dijkstra, 1959), but their search frontiers grow quickly on large networks and become memory- and time-intensive.
-Graph partitioning methods (e.g., Kernighan-Lin, multilevel partitioning/METIS, spectral cuts) reduce problem size by cutting the graph into parts (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Shi and Malik, 2000), yet cuts chosen without regard to a specific query $(s,t)$ can intersect narrow passages that many feasible sstt routes must use, which can exclude high-quality routes or require costly boundary repair.
-Bidirectional search reduces depth by growing frontiers from $s$ and $t$ (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997; Holte et al., 2016), but still explores a single coupled state space and may expand many partial routes on large graphs.
-For completeness, we note that while feasibility for unconstrained shortest paths is trivial on connected graphs, constrained variants (e.g., resource- or budget-limited paths) are NP-hard; our focus here is the unconstrained case with nonnegative weights, which admits an optimal solution via Dijkstra’s algorithm (Dijkstra, 1959).
-We propose SPHERE, a query-aware partition that aligns computation to the sstt geometry.
-Let $B_{s}(r)$ and $B_{t}(r)$ denote hop-distance spheres around $s$ and $t$.
-SPHERE starts from a radius tuple $(r_{s},r_{t})$ with guaranteed overlap $B_{s}(r_{s})\cap B_{t}(r_{t})\neq\emptyset$, then applies a monotone decrement rule that reduces the radii (when needed, both simultanously) so that $|r_{s}-r_{t}|\leq 1$ while preserving overlap, and selects the minimal balanced pair, i.e.
-no further admissible unit decrement of either radius (still keeping $|r_{s}-r_{t}|\leq 1$) leaves the intersection nonempty.
-Then it selects an anchor $a\in B_{s}(r_{s})\cap B_{t}(r_{t})$ and splits the query into $s\!\to\!a$ and $a\!\to\!t$, solved on the corresponding induced subgraphs under a tunable size cap.
-If a subproblem still exceeds the cap, the procedure recurses on its induced subgraph.
-Because the cut lies inside the overlap, concatenating subpaths at $a$ yields a valid sstt route without boundary repair.
-The partition is solver-agnostic and the independent subproblems can be solved in parallel.
-Our contributions are:
+Point-to-point routing in large weighted, undirected graphs is a canonical problem in operations research, one that emerged alongside the field’s inception.
+Exact label-setting or label-correcting methods (e.g., Dijkstra) can certify optimality [Dijkstra1959], but their search frontiers grow quickly on large networks and become memory- and time-intensive.
+Graph partitioning methods (e.g., Kernighan-Lin, multilevel partitioning/METIS, spectral cuts) reduce problem size by cutting the graph into parts [KernighanLin1970, KarypisKumar1998, ShiMalik2000], yet cuts chosen without regard to a specific query $(s,t)$ can intersect narrow passages that many feasible sstt routes must use, which can exclude high-quality routes or require costly boundary repair.
+Bidirectional search reduces depth by growing frontiers from $s$ and $t$ [Nicholson1966, Pohl1969, KaindlKainz1997, Holte2016], but still explores a single coupled state space and may expand many partial routes on large graphs.
+For completeness, we note that while feasibility for unconstrained shortest paths is trivial on connected graphs, constrained variants (e.g., resource- or budget-limited paths) are NP-hard; our focus here is the unconstrained case with nonnegative weights, which admits an optimal solution via Dijkstra’s algorithm [Dijkstra1959].
+We propose SPHERE, a query-aware source-target partitioning heuristic that aligns computation directly with the sstt geometry.
+Let $S_{R}(s)$ and $S_{R}(t)$ denote hop-distance spheres centered around $s$ and $t$.
+SPHERE initializes a radius tuple $(R_{s},R_{t})$ large enough to guarantee a nonempty overlap; $S_{R_{s}}(s)\cap S_{R_{t}}(t)\neq\emptyset$.
+SPHERE then applies a monotone decrement rule to systematically shrink the radii, maintaining a balanced scale ($|R_{s}-R_{t}|\leq 1$) until reaching the minimal balanced pair which refers to the situation where any further valid reduction of either radius would leave the intersection empty.
+From this minimal overlap, the algorithm selects an anchor node $a\in S_{R_{s}}(s)\cap S_{R_{t}}(t)$.
+This anchor effectively splits the original source-target pair sstt into two independent subproblems, $s\to a$ and $a\to t$, which are then solved on their respective induced subgraphs subject to a tunable bound.
+This tunable bound is a bound that limits the maximum hop-radius of a subproblem which we refer to as $R_{\max}$.
+If a resulting subproblem still exceeds this cap, the procedure simply recurses on its induced subgraph.
+Crucially, because the anchor $a$ sits inside the shared overlap, it exists in both induced subgraphs simultaneously.
+This guarantees that the local $s\to a$ path and the local $a\to t$ path will physically connect at $a$, seamlessly forming a continuous, valid sstt route without requiring any secondary algorithms to repair broken links across the cut.
+Finally, SPHERE is solver-agnostic and generates independent subproblems that can be executed in parallel.
+We make the following contributions:
   1. 1.
-    Query-aware partition.
-We introduce a simple heuristic that adaptively splits the graph based on the geometry of the source-target pair, ensuring that each part preserves connectivity and can be rejoined without additional repair.
+    Query-Aware Geometric Partitioning: We introduce a heuristic that adaptively partitions the graph based on the specific sts-t geometry by cutting inside the last nonempty overlap of hop-distance spheres.
+This ensures that each part preserves local connectivity and can be rejoined at an anchor node without the need for costly boundary repair.
   2. 2.
-    Parallel subinstances.
-The subproblems generated by SPHERE are independent induced subgraphs, naturally exposing parallelism and enabling scalable routing on distributed systems with effort scaling to hardware.
+    Theoretical Guarantees: We formalize the spherical partitioning algorithm and prove that our cut placement intrinsically preserves global route feasibility.
+Furthermore, for unweighted, undirected graphs we establish theoretical conditions under which the recursive application of this partition guarantees a globally optimal shortest path.
   3. 3.
+    Solver-Agnostic Parallel Execution: The subproblems generated by SPHERE are completely independent induced subgraphs.
+This allows the subproblems to be solved in parallel.
+Because SPHERE generates standard graphs, start nodes, and anchor nodes to the downstream solver, it is completely solver-agnostic.
+Any off-the-shelf routing algorithm (e.g.
+Dijkstra) can be used to compute local paths without the need to adjust the solver’s underlying structure.
+  4. 4.
     Empirical validation.
-On large road graphs, SPHERE achieves faster runtimes than Louvain-based routing and a METIS-based pipeline while matching or improving route quality, benefiting further from inherent parallelism.
-SPHERE is complementary to downstream solvers and can wrap Dijkstra or A* (Dijkstra, 1959; Hart et al., 1968), as well as modern speedup indices such as ALT/landmarks (Goldberg and Harrelson, 2005), reach-based pruning (Gutman, 2004), contraction hierarchies (Geisberger et al., 2008), hub labeling (Abraham et al., 2011), and transit-node routing (Bast et al., 2016); we evaluate with common solvers and baselines and discuss integration in Section 5.
-Next, we motivate these design choices by examining where existing approaches trade-off feasibility, and instance size, before positioning SPHERE within the literature and formalizing our approach.
+On large-scale networks with over a million nodes and edges, SPHERE achieves faster runtimes and smaller optimality gaps than other state-of-the-art sstt routing approaches.
+Crucially, it also substantially mitigates heavy-tail runtime slowdowns observed in standard methods, resulting in highly stable and predictable execution times across varying queries.
 2 Motivation
-Routing at scale must meet three criteria simultaneously: control per-instance complexity, make an $s$-$t$ route feasible, and remain compatible with trusted downstream solvers while enabling parallel execution.
-Existing tools typically sacrifice at least one of these: Labeling methods (Feillet et al., 2004; Irnich and Desaulniers, 2005) preserve feasibility but generate a combinatorial explosion of partial routes on large graphs.
-Graph partitioners (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Shi and Malik, 2000) control instance size, but because their boundaries are chosen without regard to $s$ and $t$, they may cut across a bottleneck that every feasible sstt path must traverse, risking the loss of valid routes.
-Bidirectional search (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997; Holte et al., 2016) reduces search depth by expanding from both ends, $s$ and $t$, but the two frontiers still form a single combined search space and therefore cannot be partitioned into smaller, independent pieces suitable for parallel solution.
-SPHERE overcomes these trade-offs by aligning the partition with the specific request.
-To ensure controllable size, it expands hop-distance spheres from $s$ and $t$ until reaching the last nonempty overlap, where a user-defined radius limits the diameter of each subproblem.
-Second, for feasibility, SPHERE selects an anchor $a$ in the overlap and solves $s\!\to\!a$ and $a\!\to\!t$ on their respective induced subgraphs, so cuts always lie inside the overlap and concatenation at $a$ is guaranteed to generate a feasible sts-t route.
-This preserves all local connectivity within each part and ensures that every feasible subpath remains feasible in the global graph.
-Third, the resulting subproblems are independent, enabling parallel execution across multiple cores, and the design is solver-agnostic which means that any downstream optimizer, from exact Mixed-Integer Programming (MIP) on small pieces to strong metaheuristics on larger ones (Pisinger and Ropke, 2007; Vidal, 2022), can be applied without modification.
+Algorithms for routing at scale inherently struggle to balance search space complexity, route exactness, and practical scalability [bast2016route].
+Building on these established challenges, routing at scale must meet three criteria simultaneously: control per-instance complexity, guarantee route feasibility, and enable parallel execution with standard solvers.
+Existing tools typically sacrifice at least one of these:
+Labeling methods preserve feasibility, but generate a combinatorial explosion of partial routes due to the complexity of the topology of large graphs [Feillet2004, IrnichDesaulniers2005].
+Graph partitioners such as [KernighanLin1970, KarypisKumar1998, ShiMalik2000] successfully control instance size, but their static boundaries ignore the specific $s$ and $t$ locations.
+This risks cutting across critical bottlenecks and losing feasible routes.
+Bidirectional search [Nicholson1966, Pohl1969, KaindlKainz1997, Holte2016] reduces depth but explores a single coupled state space that cannot be partitioned into independent pieces, making it difficult to enforce parallelism.
+While static partitioning methods like Corridor routing could theoretically be parallelized, they often require shared boundary information or synchronization across a global quotient graph to ensure path consistency.
+In contrast, SPHERE’s subproblems are mathematically independent by construction.
+Because the anchor node $a$ is guaranteed to exist in both induced subgraphs, no inter-process communication or boundary repair is required during the solve phase, allowing for true solver-agnostic parallel execution.
+In comparison, SPHERE avoids combinatorial explosion by bounding subproblems within hop-distance spheres, strictly limited by a tunable bound ($R_{\max}$).
+Additionally, SPHERE cuts exclusively inside the shared sstt overlap and preserves local connectivity and guarantees a feasible route upon concatenation without any need for boundary repair.
+The induced subgraphs generated by SPHERE naturally expose parallel execution and act as a solver-agnostic framework, allowing practitioners to drop-in any trusted downstream optimizer, from exact Mixed-Integer Programming (MIP) to strong metaheuristics [PisingerRopke2007, Vidal2022], without modification.
 In short, SPHERE offers a partition strategy that controls complexity, preserves feasibility, and exposes natural parallelism while remaining compatible with the solvers practitioners already trust.
@@
 Point-to-point routing on large weighted graphs has been studied from several perspectives.
-Classical exact methods such as label-setting and label-correcting algorithms (e.g., Dijkstra) can guarantee global optimality on connected graphs with nonnegative weights (Dijkstra, 1959).
-Dynamic-programming and labeling variants extend this to richer settings and can certify optimality, but they often face an explosion in the number of labels on large graphs (Feillet et al., 2004; Irnich and Desaulniers, 2005).
-Approximation schemes exist for restricted variants (Hassin, 1992), trading accuracy for speed, yet these techniques operate on the full graph and therefore do not reduce the instance size before optimization.
-Graph partitioning offers a complementary idea: simplify the problem by splitting the graph.
-Kernighan-Lin improves an initial cut by greedy node swaps to reduce the number of edges crossing between parts while keeping parts balanced (Kernighan and Lin, 1970).
-Multilevel schemes such as METIS repeatedly coarsen the graph, compute a cut on the coarse instance, and refine it during uncoarsening, which gives strong scalability in practice (Karypis and Kumar, 1998).
-In routing pipelines, these partitions are often used to build a quotient graph for a coarse path that is later refined within parts.
-A limitation of all such methods is that the cut is chosen without regard to the specific start-target pair $(s,t)$, so the boundary can lie across a narrow passage that every feasible $s$-$t$ path must use.
-Once this happens, the coarse or restricted routing stage may miss or degrade the necessary connection.
-Bidirectional search reduces search depth by expanding frontiers from both ends and meeting near the middle (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997).
-The MM algorithm of Holte et al.
-(2016) provides a priority rule that ensures expansions do not go beyond the midpoint of the optimal solution.
+Classical exact methods such as label-setting and label-correcting algorithms (e.g., Dijkstra) can guarantee global optimality on connected graphs with nonnegative weights [Dijkstra1959].
+Dynamic-programming and labeling variants extend this to richer settings and can certify optimality, but they often face an explosion in the number of variables on large graphs [Feillet2004, IrnichDesaulniers2005].
+Approximation schemes [Hassin1992] exist, trading accuracy for speed, yet these techniques operate on the full graph and therefore do not reduce the instance size before the optimization.
+3.1 Graph Partitioning and Parallel Solvers
+Graph partitioning offers a complementary idea: simplify the problem by splitting the graph, a strategy specifically suited for large-scale routing.
+Kernighan-Lin improves an initial cut by greedy node swaps to reduce the number of edges crossing between parts while keeping parts balanced [KernighanLin1970].
+Multilevel schemes such as METIS repeatedly coarsen the graph, compute a cut on the coarse instance, and refine it during uncoarsening, which gives strong scalability in practice [KarypisKumar1998].
+In routing pipelines, these partitions, along with community-based heuristics like the Louvain method, are often used to build a quotient graph for a coarse path or corridor that is later refined within parts [Blondel2008].
+A limitation of all such methods is that the cut is chosen without regard to the specific start-target pair $(s,t)$, so the boundary can lie across a narrow passage that every feasible sstt path must use.
+Once this happens, the coarse or restricted routing stage may miss or degrade the necessary connection leading to potentially infeasible routes.
+Parallel computation for shortest-path problems has also been studied to accelerate routing.
+For instance, Rozhon2023 analyze parallel Breadth-First Search and approximate distances within subgraphs.
+However, this approach is mainly theoretically analyzed.
+3.2 Machine Learning and Deep Learning-Based Approaches
+Recent advancements in neural combinatorial optimization have introduced Deep Learning models capable of solving complex routing tasks.
+Notable examples include Policy Optimization with Multiple Optima (POMO) [Kwon2020] and Multi-Task Vehicle Routing Solver with Mixture-of-Experts (MvMOE) [Zhou2024].
+These neural solvers are primarily designed for distribution-trained variants of the Traveling Salesperson Problem (TSP) and Vehicle Routing Problem (VRP).
+Those methods operate effectively on small to medium-scale graphs.
+3.3 Bidirectional Search
+Bidirectional search reduces search depth by expanding frontiers from both ends and meeting near the middle [Nicholson1966, Pohl1969, KaindlKainz1997].
+The MM algorithm of Holte2016 provides a priority rule that ensures expansions do not go beyond the midpoint of the optimal solution.
 These methods improve exploration efficiency but still operate within a single coupled search space and do not produce small, independent instances.
@@
 4.1 Spherical subgraphs
-We begin by fixing distance and the local “window” around a node.
-In the following, let every graph $G=(V,E)$ be a finite, connected, undirected graph with positive edge weights.
-Unless stated otherwise we measure distance in hops (unweighted shortest path); all statements below remain valid for the weighted shortest-path distance.
-Definition 1 (Spheres).
+We begin by formalizing the local "window" around a node using hop-based distance.
+Throughout this work, let $G=(V,E)$ denote a finite, connected, and undirected graph with positive edge weights.
+Unless otherwise specified, we measure distance in hops (unweighted shortest path), though our formulations remain valid for weighted shortest-path distances.
+For a comprehensive summary of the mathematical notation used in this paper, please refer to Appendix A.1.
+Definition 4.1 (Spheres)
 For $R\in\mathbb{N}$ and a reference node $\tilde{v}\in V$, the (closed) sphere is
-$S_{R}(\tilde{v})\;:=\;\{\,v\in V:\ d_{G}(\tilde{v},v)\leq R\,\}\subseteq V,$
+$S_{R}(\tilde{v})\;\coloneqq\;\{\,v\in V:\ d_{G}(\tilde{v},v)\leq R\,\}\subseteq V,$
 where $d_{G}(\tilde{v},v)$ is the unweighted shortest path distance between $\tilde{v}$ and $v$.
 […]
-Definition 2 (Spherical subgraph).
+Definition 4.2 (Spherical subgraph)
 […]
-$E^{\mathrm{sp}}_{R}(\tilde{v})=\bigcup_{v\in S_{R}(\tilde{v})}\mathcal{E}\!\bigl(\mathrm{SP}(\tilde{v},v)\bigr),$
+$E^{\mathrm{sp}}_{R}(\tilde{v})\coloneqq\Bigl\{\{i,j\}\in E:$ $d_{G}(\tilde{v},i)\leq R,\;d_{G}(\tilde{v},j)\leq R,$
+$\bigl|d_{G}(\tilde{v},i)-d_{G}(\tilde{v},j)\bigr|=1\Bigr\}.$
 […]
-Definition 3 (Induced spherical subgraph).
+Definition 4.3 (Induced spherical subgraph)
+[…]
 […]
@@
 4.2 Partition cut
-With these building blocks in place, we now align the local windows to a specific query $(s,t)$. For radii $(R_{s},R_{t})\in\mathbb{N}^{2}$, define the overlap
-$O(R_{s},R_{t})\;:=\;S_{R_{s}}(s)\cap S_{R_{t}}(t).$
-We control how radii evolve, how they are initialized, and how an anchor is chosen in the overlap.
-Definition 4 (Decrement rule).
-A decrement rule is a map $\mathrm{decru}:\mathbb{N}^{2}\to\mathbb{N}^{2}$ that is monotone in the lexicographic order, i.e.,
+With these geometric building blocks established, we now align these local windows to a specific query $(s,t)$.
+Definition 4.4
+For radii $(R_{s},R_{t})\in\mathbb{N}^{2}$, define the overlap as
+$O(R_{s},R_{t})\;\coloneqq\;S_{R_{s}}(s)\cap S_{R_{t}}(t).$
+In order to identify a precise splitting point, we control how the radii evolve from their initial state to a minimal configuration.
+This necessitates defining the properties of decrement rules, which systematically shrink the search space.
+Definition 4.5 (Decrement rule)
+A decrement rule
+$\mathrm{decru}:\mathbb{N}^{2}\to\mathbb{N}^{2}$
+is a map that is monotone in the lexicographic order, i.e.,
 $(x_{1},x_{2})\leq_{\mathrm{lex}}(y_{1},y_{2})\;\Longleftrightarrow$ $\begin{aligned} &x_{1}<y_{1},\\ &\text{or }(x_{1}=y_{1}\text{ and }x_{2}\leq y_{2}).\end{aligned}$
 and never increases any coordinate.
-Example (monotone decrement).
-A simple rule that satisfies the definition is
+Example 4.6 (monotone decrement)
+A standard rule that satisfies these properties is one that prioritizes shrinking the larger sphere to keep the search balanced:
 $\mathrm{decru}(R_{s},R_{t})=\begin{cases}(R_{s}{-}1,R_{t}),&R_{s}\geq R_{t},\\[2.0pt] (R_{s},R_{t}{-}1),&\text{otherwise.}\end{cases}$
-It is coordinatewise nonincreasing, and it is lex-monotone: if $(R_{s},R_{t})\leq_{\mathrm{lex}}(R^{\prime}_{s},R^{\prime}_{t})$, then either $R_{s}<R^{\prime}_{s}$ (so the first branch cannot produce a larger first coordinate) or $R_{s}=R^{\prime}_{s}$ with $R_{t}\leq R^{\prime}_{t}$, in which case both images either decrement the same coordinate or the smaller pair cannot “overtake” the larger one in lexicographic order.
-The map $\mathrm{decru}$ from Definition 4 determines how radii shrink once chosen.
-To make the procedure operative, we also need: (i) an initial choice of radii that guarantees the spheres overlap at the start, and (ii) a way to convert any nonempty overlap into a concrete split point.
-These are captured by the following two rules.
-Definition 5 (Starting rule).
-A starting rule $\mathrm{staru}:V\times V\to\mathbb{N}^{2}$ returns $(R_{s}^{0},R_{t}^{0})$ with nonempty overlap $O(R_{s}^{0},R_{t}^{0})\neq\emptyset$ (e.g., $R_{s}^{0}{+}R_{t}^{0}\geq d_{G}(s,t)$; a stronger but safe choice is $S_{R_{s}^{0}}(s)\cup S_{R_{t}^{0}}(t)=V$).
-Definition 6 (Anchor rule).
-An anchor rule $\mathrm{anru}:\mathcal{P}(V)\to V$ selects $a\in O(R_{s},R_{t})$ (e.g., uniform random).
-Bringing these pieces together, the three rules: starting rule for initialization (Definition 5), decrement rule for shrinking (Definition 4), and anchor rule for selecting the split point (Definition 6), jointly define the following algorithm, which we refer to as the Partition cut:
+This rule ensures the radii remain balanced ($|R_{s}-R_{t}|\leq 1$) throughout the process.
+If $(R_{s},R_{t})\leq_{\mathrm{lex}}(R^{\prime}_{s},R^{\prime}_{t})$, the smaller pair cannot "overtake" the larger one, preserving the order of operations.
+We make this procedure operational by defining two auxiliary rules for initialization and anchor selection:
+Definition 4.7 (Starting rule)
+A starting rule
+$\mathrm{staru}:V\times V\to\mathbb{N}^{2}$
+returns $(R_{s}^{0},R_{t}^{0})$ with nonempty overlap $O(R_{s}^{0},R_{t}^{0})\neq\emptyset$ (e.g., $R_{s}^{0}{+}R_{t}^{0}\geq d_{G}(s,t)$; a stronger but safe choice is $S_{R_{s}^{0}}(s)\cup S_{R_{t}^{0}}(t)=V$).
+Definition 4.8 (Anchor rule)
+An anchor rule
+$\mathrm{anru}:\mathcal{P}(V)\to V$
+selects $a\in O(R_{s},R_{t})$ (e.g., uniform random).
+The starting rule for initialization (Definition 4.7), decrement rule for shrinking (Definition 4.5) and anchor rule for selecting the split point (Definition 4.8), together define Algorithm 1, which we refer to as the Partition cut.
 Algorithm 1 Partition cut
@@
 10:return $(\bar{R}_{s},\bar{R}_{t},O,a)$
-The intuition behind Algorithm 1 can be explained as follows and is visualized in Figure 2.
-Grow spheres around $s$ and $t$ until they overlap, then shrink via $\mathrm{decru}$ until the next shrink would empty the overlap.
-Pick an anchor $a$ according to Definition 6 in this last nonempty overlap, and split the query into $s\!\to\!a$ and $a\!\to\!t$ subproblems on the corresponding induced spheres.
+The core intuition behind the Partition Cut (visualized in Figure 2) is to identify the smallest possible geometric region where the influence of $s$ and $t$ meet.
+Initially, spheres are grown around the terminals until they intersect.
+We then apply the decrement rule, $\mathrm{decru}$, to systematically shrink these radii until any further reduction would result in an empty intersection.
+By selecting an anchor node $a$ within this final, minimal overlap (Definition 4.8), we effectively decompose the original $s\to t$ query into two independent $s\to a$ and $a\to t$ subproblems.
+These subproblems are then solved within their respective induced spherical subgraphs.
 […]
-Figure 2: Last nonempty overlap. Blue and red spheres overlap in the green region containing both $a$ and $a^{\prime}$. Potential anchor node $a$ is selected according to Definition 6.
+Figure 2: Last nonempty overlap. Blue and red spheres overlap in the green region containing both $a$ and $a^{\prime}$. Potential anchor node $a$ is selected according to Definition 4.8.
 A natural question is whether the Partition cut of Algorithm 1 actually preserves feasibility of the original $s{\to}t$ routing problem.
-In other words, does splicing the two subpaths at the chosen anchor still yield a valid global route?
+In other words: Does splicing the two subpaths at the chosen anchor still yield a valid global route?
 The following statement makes this precise:
-Proposition 4.1 (Feasibility preservation).
+Proposition 4.9 (Feasibility preservation)
 Let $(\bar{R}_{s},\bar{R}_{t},O,a)$ be returned by Algorithm 1.
 If $P_{s\to a}$ is a feasible path in $G^{\mathrm{ind}}_{\bar{R}_{s}}(s)$ and $P_{a\to t}$ is a feasible path in $G^{\mathrm{ind}}_{\bar{R}_{t}}(t)$, then $P_{s\to a}\!\circ\!P_{a\to t}$ is a feasible $s\to t$ path in $G$.
-Proof sketch.
-Both subgraphs are induced, hence every used edge is in $G$. Since $a\in O(\bar{R}_{s},\bar{R}_{t})$, concatenating at $a$ yields a valid $s\to t$ path in $G$. ∎
+Proof 4.10.
+We provide a short proof sketch: Both subgraphs are induced, hence every used edge is in $G$. Since $a\in O(\bar{R}_{s},\bar{R}_{t})$, concatenating at $a$ yields a valid $s\to t$ path in $G$.
 4.3 Recursive spherical partitioning
-The following algorithm extends the partition cut from Algorithm 1 into a recursive partitioning routine.
-Its goal is to partition a global $s\!\to\!t$ routing task into a set of smaller subproblems, each confined to an induced sphere whose radius does not exceed a user-defined bound $R_{\max}$.
-These subproblems are returned as task triples $(H,u,w)$, meaning: “route from $u$ to $w$ within the induced subgraph HH.”
+The following procedure extends the partition cut of Algorithm 1 into a recursive framework.
+The objective is to decompose a global $s\to t$ routing task into a collection of manageable subproblems, each strictly confined to an induced spherical subgraph where the radius does not exceed a bound $R_{\max}$.
+For a comprehensive overview of all configurable parameters, please refer to Appendix A.2.
+Each subproblem is formalized as a task triple $(H,u,w)$, representing a localized search from node $u$ to node $w$ within the induced subgraph $H$.
+We treat the resulting task set $\mathcal{T}$ as a sequence ordered along the $s\to t$ trajectory.
+This ordering ensures that concatenating the locally optimal paths induces a valid global path.
+For simplicity, we refer to this locally optimal path as the path returned by Algorithm 2.
 Algorithm 2 Recursive spherical partitioning
 1graph $G=(V,E)$, terminals $(s,t)$,
-2     radius cap $R_{\max}$, rules $\mathrm{staru},\mathrm{decru},\mathrm{anru}$
+2     bound $R_{\max}$, rules $\mathrm{staru},\mathrm{decru},\mathrm{anru}$
 3  set $\mathcal{T}$ of task triples $(H,u,w)$
@@
 (G,s,t)(G,s,t)Partition cut $\Rightarrow$ $(R_{s},R_{t},O,a)$$G^{\mathrm{ind}}_{R_{s}}(s),\ s\to a$$G^{\mathrm{ind}}_{R_{t}}(t),\ a\to t$if $R_{s}\leq R_{\max}$: record; else: recurseif $R_{t}\leq R_{\max}$: record; else: recurse
-Figure 3: Recursive splitting. Any side above the radius cap is split again by Algorithm 1; leaves are independent subproblems.
+Figure 3: Recursive splitting. Any side above the bound is split again by Algorithm 1; leaves are independent subproblems.
 Figure 3 illustrates this recursive structure.
 Each partition cut splits the problem into two boxes, one for the start-side and one for the target-side.
-If the radius on a side is within the threshold, the box is recorded as a leaf; otherwise, it is expanded by another partition cut.
+If the radius on a side is within the bound, the box is recorded as a leaf; otherwise, it is expanded by another partition cut.
 4.3.1 Baseline application
-The recursive partition in Algorithm 2 produces a set of leaf tasks, each represented by a triple $(H,u,w)$ where $H$ is an induced spherical subgraph and $u,w$ are its boundary terminals.
-At this stage, the global $s\to t$ routing problem has been reduced to a collection of independent local subproblems, each small enough to be solved efficiently by standard methods.
+The recursive decomposition defined in Algorithm 2 yields a set of leaf tasks, each formalized as $(H,u,w)$.
+This process effectively reduces the high-complexity global $s\to t$ routing problem into a collection of independent, localized subproblems.
+Because these subproblems are spatially confined and bounded by the bound $R_{\max}$, they are sufficiently small to be solved efficiently using standard routing methods or off-the-shelf optimizers.
 Algorithm 3 Baseline application with spherical partitioning
 1:graph $G=(V,E)$, terminals $(s,t)$,
-2:     radius cap $R_{\max}$, solver $\mathrm{optimize}(\cdot)$
+2:     bound $R_{\max}$, solver $\mathrm{optimize}(\cdot)$
 3:feasible route $\mathcal{R}$
 […]
-5:for all $(H,u,w)\in\mathcal{T}$ do
+5:for each $(H,u,w)\in\mathcal{T}$ do
 6:  $\mathcal{R}\leftarrow\mathcal{R}\circ\mathrm{optimize}(H,u,w)$
@@
 8:return $\mathcal{R}$
-Algorithm 3 describes how to turn the recursive spherical partitioning into a full routing solution.
-First, the input graph $G$ with terminals $s,t$ is partitioned into a collection of smaller subproblems (line 3).
-Each subproblem $(H,u,w)$ is an induced subgraph $H$ together with its local terminals $u$ and $w$.
-Then, for each subproblem, an arbitrary downstream solver $\mathrm{optimize}(H,u,w)$ is applied to compute a local path (lines 4-5).
-These local solutions are concatenated in the order determined by the anchors of the partition, resulting in a valid global route $\mathcal{R}$ from $s$ to $t$ (line 6).
-Since the subproblems are independent, they can also be solved in parallel, and the procedure remains agnostic to the choice of solver.
+Algorithm 3 details the integration of recursive spherical partitioning into a complete routing pipeline.
+The procedure begins by decomposing the input graph $G$ and terminals $(s,t)$ into a sequence of smaller, manageable subproblems with resulting task $(H,u,w)$.
+Subsequently, an arbitrary downstream solver, $\mathrm{optimize}(H,u,w)$, is applied to each subproblem to compute a local path.
+These local solutions are then concatenated in the order established by the partition’s anchors, forming a globally valid $s\to t$ route $\mathcal{R}$.
+A critical advantage of this framework is that because the subproblems are mutually independent, they can be solved in parallel.
+Furthermore, the procedure remains strictly solver-agnostic, allowing the utilization of various downstream optimizer without modification to the underlying partitioning logic.
+4.4 Optimality Conditions
+[…]
+In the following, we demonstrate that Algorithm 1, when applied to unweighted graphs using the monotone decrement rule from Example 4.6, preserves optimality.
+We thus provide a stronger result compared to the the general weighted setting, where Proposition 4.9 ensures the preservation of route feasibility only.
+Furthermore, this result allows to compute a globally shortest path without invoking any additional optimization routines or solver-specific procedures, provided we apply the recursion until the radii are equal to $1$.
+We begin by formalizing the Strong Minimality Condition, which characterizes the minimal non-empty overlap induced by the monotone decrement rule in terms of the associated radius tuple $(R_{s},R_{t})$.
+Definition 4.11 (Strong Minimality Condition).
+A tuple $(R_{s},R_{t})$ is strongly minimal if
+$O(R_{s},R_{t})\neq\emptyset,$
+$O(R_{s}-1,R_{t})=\emptyset,$
+$O(R_{s},R_{t}-1)=\emptyset.$
+The next result shows by combining Definitions 4.4 and 4.11 that strong minimality forces all overlap vertices to lie simultaneously on the boundaries of both spheres determined by $(R_{s},R_{t})$.
+Lemma 4.12 (Boundary property).
+If $(R_{s},R_{t})$ is strongly minimal, then for every $v\in O(R_{s},R_{t})$,
+$d(s,v)=R_{s}\quad\text{and}\quad d(t,v)=R_{t}.$
+Proof 4.13.
+Let $v\in O(R_{s},R_{t})$.
+If $d(s,v)\leq R_{s}-1$, then $v\in O(R_{s}-1,R_{t})$, contradicting strong minimality.
+Hence $d(s,v)=R_{s}$.
+The argumentation for $t$ is analogous, yielding $d(t,v)=R_{t}$.
+Having established the boundary property of strongly minimal overlaps, we now derive the key consequence of the monotone decrement rule.
+The following Meet-in-the-middle equality formalizes that a strongly minimal radius tuple satisfies
+$R_{s}+R_{t}=d(s,t),$
+thereby identifying the last overlap of two spheres.
+Lemma 4.14 (Meet-in-the-middle equality).
+If $(R_{s},R_{t})$ is strongly minimal, then
+$R_{s}+R_{t}=d(s,t).$
+Proof 4.15.
+Choose $a\in O(R_{s},R_{t})$. Since $a\in S_{R_{s}}(s)$ and $a\in S_{R_{t}}(t)$, there exists an sstt path via $a$ of length at most $R_{s}+R_{t}$, hence $d(s,t)\leq R_{s}+R_{t}$.
+Assume for contradiction that $d(s,t)\leq R_{s}+R_{t}-1$.
+Let $P$ be a shortest sstt path of length $L=d(s,t)$.
+Let $v$ be the vertex on $P$ at distance $R_{s}$ from $s$.
+Since $P$ is the shortest sstt path, then $d(s,v)=R_{s}$ and $d(t,v)=L-R_{s}\leq R_{t}-1$.
+Thus $v\in O(R_{s},R_{t}-1)$ contradicts strong minimality.
+Therefore $d(s,t)\geq R_{s}+R_{t}$, and equality follows.
+Lemma 4.14 constitutes the central argument of our main optimality theorem: it ensures that any anchor selected from the minimal overlap lies on a globally shortest path and thus enables recursive preservation of optimality.
+Note that all definitions and lemmas apply to any terminal pair $(u,w)$ (replacing $(s,t)$), in every recursive subproblem.
+Theorem 4.16 (Global preservation).
+Let $G=(V,E)$ be a finite, connected, undirected, and unweighted graph and assume that the monotone decrement rule is used.
+Then the concatenated sstt path returned by Algorithm 2 is a globally shortest path in $G$.
+Proof 4.17.
+We argue by induction over the recursion tree:
+Local optimal split.
+Fix any subproblem with endpoints $(u,w)$ and a strongly minimal pair $(R_{u},R_{w})$ as induced by undirected graphs and the monotone decrement rule with anchor $a\in O(R_{u},R_{w})$.
+Applying Lemma 4.12, we observe that the anchor $a$ lies on the boundaries of both spheres, such that $d(u,a)=R_{u}$ and $d(a,w)=R_{w}$.
+It follows from Lemma 4.14 that these radii sum exactly to the distance between $u$ and $w$, i.e.
+$R_{u}+R_{w}=d(u,w)$.
+Hence
+$d(u,a)+d(a,w)=d(u,w),$
+so $a$ lies on a shortest uuww path and the split $(u,w)\mapsto(u,a),(a,w)$ is consistent with global optimality for that subproblem.
+Inductive step.
+Assume recursively computed solutions are shortest for the two child subproblems $(u,a)$ and $(a,w)$.
+Their concatenation has length $d(u,a)+d(a,w)=d(u,w)$, hence is shortest for the parent subproblem.
+Applying this from leaves to the root yields a globally shortest sstt path.
 5 Numerical Experiments
@@
 5.1 Dataset
-For our empirical evaluation we use the well-known West-USA-road benchmark suite from the 9th DIMACS Implementation Challenge on Shortest Paths (DIM, 2006).
-These graphs encode real-world road networks at the level of individual intersections and road segments.
-Each node corresponds to a road intersection or endpoint, and each edge represents a street segment with a positive weight equal to its physical length.
-The dataset is widely used in the routing literature as a benchmark for large-scale shortest-path computations, since it combines realistic geometry, heterogeneous edge weights, and large graph sizes (millions of nodes and edges).
+For our empirical evaluation, we use the West-USA-road benchmark suite from the 9th DIMACS Implementation Challenge on Shortest Paths [DIMACS2006SP].
+This dataset represents real-world road networks.
+Nodes correspond to intersections and edges represent road segments with positive weights equal to physical length.
 5.2 Experimental Setup
-For each problem seed $p\in\{1,\dots,30\}$ we draw a source-target pair $(s,t)$ uniformly at random from the node set (rejecting $s=t$).
-For each $(s,t)$, we run five independent trials with inner seeds $q\in\{1,\dots,5\}$ to capture variability from randomized components.
-For instance, the anchors are selected randomly according to Definition 6.
-We report results per instance and aggregate by averaging across the $5$ inner seeds for each problem seed.
-Given $(s,t)$, we apply Algorithm 2 and selects an anchor node uniformly at random from this set of nodes (the randomness is controlled by the inner seed $q$).
-The graph is partitioned into subgraphs around $s$ and $t$ until a maximum radius threshold is met.
-$R_{max}$ from Algortihm 3 is set to $1800$.
-Throughout development we found a robust rule of thumb: when routing within a spherical subgraph of hop radius $R$, choosing $R_{\max}\in[\,R/2,\;R/1.7\,]$ delivers high-quality solutions at both low computational cost and effort.
-For the West-USA instances, a single sphere with $R\approx 3200$ covers the region; thus $R_{\max}=1800$ lies squarely in this band, balancing overlap likelihood and leaf size.
-Each subgraph shortest-path subproblem is solved with Dijkstra’s algorithm.
-Because our spherical partitioning produces independent induced subgraphs, we parallelize these subgraph solves and report end-to-end wall-clock time (including partition cut and final concatenation).
-By contrast, the baselines do not naturally partition into independent leaf tasks so we do not apply the same solving subproblem parallelization to them.
-Baselines.
-As baselines, we first include an optimal reference obtained by running Dijkstra’s algorithm directly on the full graph.
-To compare against partition-based approaches, we also evaluate a corridor-routing strategy in which the graph is partitioned into $k=64$ cells using PyMetis; a path is computed on the resulting quotient graph, and Dijkstra’s algorithm is then applied within the induced corridor to recover a detailed route (Karypis and Kumar, 1998).
-In addition, we test a partition heuristic based on community detection via the Louvain method: routing is first performed on the community graph, and the solution is subsequently refined by applying Dijkstra on the induced subgraph spanned by the relevant communities (default parameters; seeded where applicable) (Blondel et al., 2008).
-The corridor-routing and Louvain pipelines introduce partitions, yet can not be solved independently and not conditioned on the specific $(s,t)$, so they likewise do not yield independent leaves that one could solve with a plug-in solver and then concatenate repair-free (e.g.
-fixing feasibility) (Karypis and Kumar, 1998; Blondel et al., 2008).
-Metrics.
-Let $m\in\{\textsc{Sphere},\ \textsc{Corridor},\ \textsc{Louvain}\}$ denote the method.
+For each problem instance $p$ (with according instance seed), we draw a source-target pair $(s,t)$ uniformly at random.
+To capture variability from the randomized anchor selection, we run five independent trials with inner seeds $q$ for each source-target pair.
+We report results aggregated by the median and mean across these seeds.
+Further details regarding graph dimensions and seed selection are provided in Appendix A.3.
+We apply Algorithm 2 with a maximum radius bound of $R_{\max}=1800$ (see Appendix A.3 for tuning rationale).
+Because SPHERE generates independent induced subgraphs, we solve the resulting leaf tasks in parallel using Dijkstra’s algorithm and report the end-to-end wall-clock time, including the partitioning and concatenation overhead.
+[^1] The graph data and implementation will be made publicly available upon the manuscript’s acceptance.
+We compare SPHERE against an optimal reference using Dijkstra’s algorithm on the full graph, as well as Louvain and Corridor routing.
+Unlike SPHERE, these static partitioning schemes compute cuts globally rather than conditioning on the specific $(s,t)$ query.
+Consequently, they cannot guarantee feasibility without boundary repair, nor do they naturally yield the mathematically independent subproblems required for the same drop-in parallel execution framework applied to our method.
+While static partitions could theoretically be parallelized, they often require shared boundary synchronization or global quotient graphs to ensure path consistency, whereas SPHERE subproblems are self-contained by construction.
+(See Appendix A.3 for implementation and library details).
+Let $m\in\{\emph{SPHERE},\ \textsc{Corridor},\ \textsc{Louvain}\}$ denote the method.
 For each problem seed $p$ and inner seed $q$, let $C_{p,q}^{(m)}$ be the route cost returned by $m$, and let $C_{p}^{\star}$ be the shortest-path cost from Dijkstra’s algorithm on the full graph for the same $(s,t)$.
@@
 Across 30 West-USA instances with five inner seeds each, SPHERE achieves faster end-to-end runtimes than the three baselines and matches near-optimal route quality.
-We use performance and accuracy profiles for visualization (Dolan and Moré, 2002; Beiranvand et al., 2017).
+We use performance and accuracy profiles for visualization [dolan2002benchmarking, beiranvand2017best].
 Both are computed per instance from the median over seeds.
@@
 Figure 4: Performance and accuracy profiles of 30 instances based on median and mean over five seeds.
-Runtime.
 Figures 4(a) and 4(c) plot the fraction of instances solved within a performance ratio $\tau$ of the best method on that instance.
@@
 This suggests that SPHERE not only improves median runtime, but also avoids the heavy-tailed slowdowns observed in both Dijkstra and Louvain, resulting in tighter runtime distributions and fewer extreme outliers, which makes SPHERE’s runtimes more predictable across different instances.
-Route quality.
-Figures 4(b) and 4(d) reports the fraction of instances whose relative optimality gap is at most $\tau$.
+Figures 4(b) and 4(d) report the fraction of instances whose relative optimality gap is at most $\tau$.
 Dijkstra attains zero gap by construction.
 SPHERE’s curve dominates the partition baselines across $\tau$, yielding smaller or equal gaps on the majority of instances.
-This confirms that cutting inside the $s$-$t$ overlap and solving on induced spheres preserves high-quality routes without boundary repair.
-Pareto dominance
+This confirms that cutting inside the sstt overlap and solving on induced spheres preserves high-quality routes without boundary repair.
 Against the partition baselines (Louvain and METIS corridor), SPHERE Pareto-dominates on 26 of 30 instances in terms of average metrics.
@@
 In other words, SPHERE consistently lies on or close to the empirical speed–quality frontier, while Louvain and corridor-routing often face a trade-off between being faster but less accurate, or more accurate but substantially slower.
-Robustness and aggregation
 Repeating the analysis with means over seeds yields profiles that do not differ materially from the median-based plots, supporting robustness to the randomized anchor selection.
-Additionally, the full results in the appendix show low dispersion across seeds.
+Additionally, the full results in Appendix A.4 show low dispersion across seeds.
+The results indicate reliability across varying query geometries and suggest that spherical decomposition effectively captures the core topological features necessary for high-quality routing, regardless of internal stochasticity.
 For SPHERE, the optimality-gap standard deviation is 0 on 16 instances and is the lowest among all methods on all but one instance.
@@
 We introduced SPHERE, a routing scheme that splits an sstt query by cutting inside the last nonempty overlap of hop-spheres around $s$ and $t$.
-This placement preserves feasibility without boundary repair, confines each side to an induced subgraph, and allows the two subproblems to be solved independently with standard shortest-path routines and in parallel.
-On West-USA-road graphs, SPHERE achieves faster end-to-end times and equal or smaller optimality gaps than a Louvain-based route and a METIS-corridor pipeline on the large majority of instances; runtimes are also more predictable, as heavy-tail slowdowns seen for the baselines largely disappear.
-In contrast to bidirectional search, which still explores a single coupled state space, SPHERE yields separable leaves.
-In contrast to static partitions whose cuts ignore $(s,t)$ and can sever bottlenecks, cutting within the overlap keeps the route intact and avoids corridor repair.
-Looking ahead, the main levers are the starting rule, the decrement rule, and the anchor selection.
-Even simple choices already perform well, but data-driven or adaptive variants should improve robustness across graphs and query distributions.
-A second priority is to tune the radius cap $R_{\max}$ automatically (per instance or region) to balance overlap likelihood against subproblem size.
-These steps retain the current simplicity and parallelism while targeting further speed and stability.
+This placement preserves feasibility without boundary repair, confines each side to an induced subgraph, and allows the subproblems to be solved independently with standard shortest-path routines and in parallel.
+On million-scale road networks, SPHERE achieves faster runtimes and smaller optimality gaps than state-of-the-art static and community-based partitioning methods.
+Notably, our approach eliminates the heavy-tail runtime outliers common in exact solvers, providing high statistical stability across randomized anchor selections.
+Future work will explore data-driven anchor selection to further improve robustness across diverse graph topologies.
+Additionally, SPHERE can serve as a wrapper for complex neural combinatorial optimizers that typically struggle with large-scale instances.
+Another direction could be to automate the tuning of the radius cap $R_{\max}$ to adaptively balance computational budgets against query complexity.
+These advancements will maintain SPHERE’s simplicity and parallelism while further enhancing speed and stability.
 Appendix A Appendix
@@
 $k_{\text{anchor}}\in\mathbb{N}$ Number of anchors to select from $O$ (often $k_{\text{anchor}}=1$).
-A.3 Performance per instance
+A.3 Implementation Details and Hyperparameters
+Graph Dimensions.
+The West-USA-road network consists of $6,262,104$ nodes and $15,248,146$ edges. The large scale of this network makes it a challenging benchmark for standard exact solvers.
+Instance and Seed Selection.
+The $30$ problem instances $p$ are defined by fixed global seeds to ensure reproducibility of the $(s,t)$ pairs.
+For each $p$, 55 inner seeds $q$ control the stochasticity of the $\mathrm{anru}$ function in Algorithm 1.
+Radius bound ($R_{\max}$).
+The tunable bound $R_{\max}$ controls the maximum diameter of the subproblems generated by Algorithm 3.
+Throughout development, we established a robust rule of thumb: when routing within a spherical subgraph of hop radius $R$, choosing $R_{\max}\in[\,R/2,\;R/1.7\,]$ delivers high-quality solutions at both low computational cost and effort.
+For the West-USA benchmark instances, a single sphere with $R\approx 3200$ generally covers the necessary region.
+Thus, our chosen bound of $R_{\max}=1800$ lies squarely in this optimal band, carefully balancing the likelihood of an sstt overlap against the complexity of the resulting leaf tasks.
+Baseline Configurations.
+The baseline pipelines were implemented using standard libraries to ensure a fair comparison.
+The corridor-routing baseline utilizes PyMetis to compute the initial $k=64$ cell partition.
+The Louvain community-detection baseline was run using its default modularity-optimization parameters, seeded where applicable to ensure deterministic execution.
+It is worth noting that because these static partitioning schemes compute cuts globally rather than conditioning on the specific $(s,t)$ query, they cannot guarantee feasibility without boundary repair, nor do they yield strictly independent leaves suitable for the same parallel execution framework applied to SPHERE.
+A.4 Performance per instance
+Continued on next page.
 Table 1: WestUSA – Routing